A screen is placed 90 cm from an object. When a converging lens is moved between the screen and the object , at two different positions images can be produced on the screen. If the distance between the positions of the lens is 20 cm , find the focal length of the lens
1
Expert's answer
2020-10-27T10:12:04-0400

d = 20 cm

u + v = 90 cm

v – u = 20 cm

v = 55 cm

u = 35 cm

Lens formula:

"\frac{1}{v} - \frac{1}{u} = \frac{1}{f}"

"\frac{1}{55} - \frac{1}{-35} = \frac{1}{f}"

"\frac{1}{55} + \frac{1}{35} = \frac{1}{f}"

"\frac{35 + 55}{1925} = \frac{1}{f}"

f = 21.38 cm

Answer: 21.38 cm